# A Survey on SAR ship classification using Deep Learning

### v1

@@ 3. RQ3
    Impact of Advanced Techniques: Techniques like pyramid features, feature fusion, and transfer learning were employed to bolster ship classification performance on SAR images [40, 64, 58, 59, 46, 25]. Handcrafted features and pyramid features demonstrated particularly promising results. However, the research in this area remains in its nascent stages, warranting further exploration of these advanced techniques.
@@ 4. RQ4
    Commonly Employed Network Architectures: Specific CNN architectures, particularly deeper models like VGG and ResNet, have shown promise in SAR ship classification due to their ability to extract complex features from SAR imagery. Interestingly, simpler and shallow CNN models, when combined with effective image processing techniques like noise reduction and image enhancement, have also achieved commendable accuracy.
@@ 5. RQ5
    Limitations and Challenges: SAR ship classification faces several critical challenges. The scarcity and imbalance of labeled SAR image data, coupled with limitations in SAR image processing techniques, hinder model performance. Additionally, relying solely on DL networks designed for optical imagery, or choosing inappropriate SAR-specific architectures, can lead to suboptimal results. Furthermore, SAR-AIS matching, while valuable, has limitations due to its inherent unreliability (e.g. lack of transmission, noisy and inconsistent messages [69]). The development of robust, automated SAR-AIS matching algorithms is crucial. Finally, SAR processing models can benefit significantly from tailored loss functions [23, 34, 48] that effectively filter out irrelevant information and focus on the features of target objects. Research on these tailored loss functions is an active area of development.
@@ 6. RQ6
    Specific Datasets and Pre-processing Techniques: Studies in the literature commonly utilize two publicly available datasets: FUSAR-Ship [17] and OpenSARShip [70]. Geometric pre-processing techniques, such as cropping, rotation, flipping, and scaling, are frequently employed to improve the generalizability of ship classification models. However, the necessity for these techniques arises from the challenges associated with small and low-resolution vessel representations in the data. Super-resolution techniques offer a potential avenue for addressing these limitations [71, 72, 73]
@@ IV Background
Several techniques are employed for SAR ship classification using DL, as illustrated in Figure 2. This section is divided into four parts. The first part explores the various DL techniques utilized in the literature for SAR ship classification. The second part examines the impact of handcrafted features on the performance of DL models. As SAR images possess unique attributes, the third part discusses how these attributes are leveraged in SAR ship classification. Finally, the fourth part investigates the use of fine-tuning for SAR ship classification.
IV-A Deep Learning Techniques
@@
As apparent from Table II, we found the origin of SAR-based ship classification using DL in 2017, when [20] classified 4 different ship types among the targets detected in 6 VV-polarized COSMO-SkyMed scenes using a simple CNN model with 2 convolution layers and 2 fully connected layers; the images were cropped to $64\times 6464 × 64$, 4 scenes were used for training, 1 for testing and 1 for validation. Their results showed that the model was able to learn the cargo ships better than the others due to their high prevalence in the training set. The advancement of SAR platforms over time has resulted in making images with better and better spatial resolutions available. [21] treats images from TerraSAR-X, with resolution of 3 meters per pixel, using CNNs. They prove that by using multiple-resolution SAR images, CNN models are good at classification tasks, achieving an f1-score of 94%. Four augmentation techniques were applied to form the training set: reflections (horizontal and vertical), rotations (-10 to 10 degrees), translation (-5 to 5 pixels) and noise addition (randomly changing pixel values to zero).
@@
[23] explores two important aspects of a DL model: attention based CNN model and a novel loss function for class-imbalanced datasets named “weighted distance measure”. A VH-polarized data set with Ground Range Detected (GRD) images is divided into 4 classes, i.e. cargo, tanker, tug and others. The results show that the model is capable of learning class representation, achieving a recognition accuracy of 84%.
@@
[27] evaluates a convolution network for detection and classification. The input data are drawn from the PALSAR-2 L-band SAR with 10 m resolution in HV polarization, and the results indicate that the CNN-based models can detect ships, but the classification performance is affected by ships that look alike.
@@
Ship classification can also be achieved using a sequence-based approach. [29] constructs a sequence input by capturing ten consecutive images of the same target at different times, degree of rotations. The authors employ a combination of CNN and long short-term memory (LSTM) with 4 categories of OpenSARShip data (boat, cargo, container, and oil tanker). Their findings suggest that the model learns improved class representations when utilizing image sequences.
@@
A CNN, when not fully connected, can be used as a feature extractor. The extracted features can then be fed to a classifier or further processed using methods like metric learning. Metric learning involves training a model to measure the similarity or distance between data points, often to cluster extracted features in a way that makes them distinguishable, thereby enhancing the model’s performance. [30] compares the performance of a CNN-only approach to a CNN combined with metric learning, on OpenSARship (bulk carrier, container, and oil tanker) focusing on data with larger sample sizes, then enhances the features extracted by the CNN by concatenating additional features learned using metric learning. The results show that the CNN combined with metric learning outperforms the simple CNN.
@@
[57] proposes a novel Hierarchical Embedding Network with Center Calibration (HENC) for few-shot ship classification. HENC utilizes a CNN to extract hierarchical features at different scales. This method achieves an accuracy of 81.03% on the six-category (cargo, tanker, tug, fishing, dredging and others) OpenSARShip dataset with few-shot training, demonstrating the effectiveness of deep models in learning class representations even with limited data.
@@
[58] discusses a deep metric learning approach that leverages a pairwise representation fusion for multi-proxy learning, and introduces a novel distribution loss to refine the sample distribution. This allows for both clustering similar classes together and training the model to make a clear distinction between all classes. The effectiveness of this three-pronged approach is evaluated on two optical ship datasets (FGSC-23 and FGSCR-42) and one SAR dataset (5-class FUSARShip).
@@
[64] introduces HDSS-Net, a novel deep learning architecture for ship classification in SAR images. This addresses challenges such as large variations in ship size, the elongated shapes of ships in SAR images, and intraclass diversity alongside interclass similarity. HDSS-Net employs a hierarchical design with three key modules. The first, a feature aggregation module (FAM), creates a multi-scale feature pyramid. The second, a Feature Boost module (FBM), utilizes rectangular convolutions to capture the elongated features of ships. Finally, a spherical space classifier (SSC) processes features in a spherical space to enhance class separability. Evaluations on OpenSARShip (three and six classes) and FUSAR-Ship (seven classes) datasets demonstrate that HDSS-Net outperforms traditional CNN methods, highlighting its potential value for maritime surveillance applications.
@@
There exists a trade-off between spatial resolution and the cost of SAR imagery. Developing algorithms that achieve accurate results using medium-resolution data is essential for broader applications. [34] uses DenseNet and metric learning to perform SAR ship classification on three-class (tanker, container, bulk carrier) and five-class (tanker, container, bulk carrier, cargo, general cargo) datasets from OpenSARShip and proposes a novel loss function to incorporate deep metric learning. To help the model learn accurate class representations, it utilizes images in triplets [74]. Furthermore, data augmentation techniques such as flipping, rotation, translation and Gaussian noise are applied to address overfitting. The results demonstrate superior performance in both the three-class and five-class categories.
@@
Each CNN layer stores important information about features, but some information may be lost when features pass from one layer to another. [33] proposes a novel multi-scale CNN (MS-CNN) method to utilize all the information from each convolutional layer. Using three classes from the OpenSARship dataset (bulk carrier, container ship and tanker), the MS-CNN, integrated with information from each convolutional layer, achieves a 5% better performance than the same CNN without integration. This improvement suggests that information loss occurs within the network.
@@
[54] uses a deep network with a novel inception-residual controller (IRC) module to improve the performance. The authors train the model on a five-category dataset from FUSARShip. The data underwent homogenization, normalization, and augmentation techniques such as flipping and rotation and, after 50 training epochs, the model with IRC achieves an accuracy of 98.71%, indicating that deeper models trained on pre-processed data can significantly improve the classification performance.
@@
[36] proposes a novel multi-scale dual-attention CNN that leverages information from convolutional layers. The authors employ a VGG-Net architecture for feature extraction and utilize a three-class dataset (cargo, container ship, and tanker) from OpenSARShip. The methodology involves modifying the VGG network to extract features from SAR images, followed by the application of squeeze-and-excitation (SE) blocks to focus on both channel-wise and positional context. SE blocks learn the importance of feature maps, promoting informative features and suppressing less useful channels. The results demonstrate the effectiveness of this method, achieving an accuracy of 92.66%. This success suggests the potential of attention mechanisms in this field. [59] proposes an enhanced vision transformer model that incorporates feature fusion modules to capture local information, followed by SE attention. The effectiveness of this approach is evaluated on a three-class OpenSARship dataset, achieving an accuracy of 76.25%.
@@
The convergence in training a DL model is controlled by several parameters, among which the learning rate is of paramount relevance. It determines how the model adjusts its internal parameters during training, and can be either fixed throughout the process or adjusted dynamically. [47] explores three learning rate management strategies and their impact on transfer learning. Findings suggest that a smooth learning rate change can improve the model’s ability to understand the data, leading to enhanced classification performance. Additionally, the authors investigate the effect of freezing different layers during transfer learning on a CNN, emphasizing the importance of layer analysis before fine-tuning. The models are first trained on the ten-category MSTAR dataset (SAR images of tanks, [75]) and then fine-tuned on a three-category OpenSARShip dataset (cargo, container ship, and bulk carrier). The results demonstrate that incorporating learning rate scheduling and thoughtful layer freezing significantly improves the model performance. This finding underscores the importance of employing effective training strategies for optimal SAR ship classification.
@@
Selecting optimal DL architectures for specific tasks can be challenging. Traditionally, researchers rely on exploring existing architectures and their applications. However, this approach often lacks a systematic methodology, leading to difficulties in finding the ideal architecture for a given classification problem. To address this challenge, [50] proposes a lightweight searched binary neural network (SBNN), obtained through network architecture search (NAS), quantized for weight reduction and incorporating patch shift processing to remove noise. The authors employ three (bulk carrier, container and tanker) and six-class (bulk carrier, container ship, tanker, cargo, general cargo and fishing) datasets from OpenSARShip, containing both VV and VH polarizations. They apply data augmentation methods such as random cropping and horizontal flipping. The SBNN achieves accuracy scores of 80.03% and 56.73% for three and six classes, respectively. NAS not only results in superior performance but also lead to a lightweight model. [53] introduces a novel mini-size searched convolutional metaformer (SCM) that leverages an updated NAS algorithm with progressive data augmentation (crop, flip, rotate, resize, cutout) to discover an efficient convolutional network baseline. This baseline is then enhanced with a transformer classifier to improve spatial awareness, and further augmented with a ConvFormer cell, combining the searched convolutional cell within a Metaformer block for superior feature extraction. Evaluated on VV- and VH-polarized datasets from OpenSARShip (3 categories, i.e. bulk carrier, container, and tanker) and FUSARShip (4 categories, i.e. bulk carrier, fishing, cargo, and tanker), SCM achieves accuracies of 82.06% and 63.9%, respectively, with a lightweight design (about half a million weights). [48] proposes a one-shot learning approach using a siamese network. The approach employs a 7-layer fully connected feature fusion that enables the network to classify 16 categories within the OpenSARShip dataset. This methodology involves three stages: (i) preprocessing the images and converting them to three dimensions by adding two additional channels (actual, filtered, and inverted SAR Image), (ii) creating image triplets containing an anchor image, a positive image, and a negative image, and (iii) training the model using a loss function that ensures the anchor image to be closer to the positive image and farther from the negative image. The results demonstrate the model’s effectiveness for ship classification.
@@
[61] addresses the challenge of handling the long-tailed distribution problem in imbalanced SAR data. The authors propose a novel architecture that incorporates a multi-branch expert network with a dual-environment sampling strategy. This network consists of three classifiers trained to handle different distribution scenarios: long-tailed, balanced, and inverse long-tailed. Intraclass dual-environment sampling adjusts sampling rates within each class based on classification confidence, creating varied distributions. Combined with center loss [76], this method enhances robustness and accuracy, outperforming existing techniques on the MSTAR and FUSAR datasets. Experiments confirmed significant improvements in recognizing both majority and minority classes, demonstrating the effectiveness of the proposed approach in balancing feature learning and improving overall and average accuracy.
@@
Due to the resolution of SAR images, research on small-sized fishing vessels has not been given sufficient importance. However, as the resolution of SAR systems increases, there is a growing need to address the classification of fishing vessels. [62] proposes a novel architecture for classifying fishing vessels, consisting of three modules: (i) the multi-feature extraction (MUL) module, which extracts features using DenseBlocks and channel-wise concatenation, (ii) the feature fusion (FF) module, which aggregates features from different layers, and (iii) the parallel channel and spatial attention (PCSA) module, which applies channel and spatial attention. The paper also proposes a method for SAR and AIS integration to create a fishing vessel dataset. This methodology was tested on three fishing vessel classes, and the model achieved an accuracy of 89.7%.

### v2

@@ 3. RQ3
    Impact of Advanced Techniques: Techniques like pyramid features, feature fusion, curriculum learning, and transfer learning were employed to bolster ship classification performance on SAR images [40, 67, 61, 62, 49, 25, 87, 88, 93]. Handcrafted features and pyramid features demonstrated particularly promising results. However, the research in this area remains in its nascent stages, warranting further investigation.
@@ 4. RQ4
    Commonly Employed Network Architectures: Specific CNN architectures, particularly deeper models such as VGG and ResNet, proved to be promising in SAR ship classification due to their ability in extracting complex features. Interestingly, simpler and shallow CNN models, when combined with effective image processing such as noise reduction and image enhancement, have also achieved commendable accuracy [55, 74].
@@ 5. RQ5
    Limitations and Challenges: SAR ship classification faces several critical challenges. The scarcity and imbalance of labeled SAR image data, coupled with limitations in SAR image processing techniques, hinder model performance. Additionally, relying solely on DL networks designed for optical imagery or choosing inappropriate SAR-specific architectures can lead to suboptimal results. Furthermore, SAR-AIS matching, while valuable, has limitations due to its inherent unreliability (e.g. lack of transmission, noisy and inconsistent messages [95]). The development of robust, automated SAR-AIS matching algorithms is crucial. Finally, SAR processing models can benefit significantly from tailored loss functions [23, 34, 51] that effectively filter out irrelevant information and focus on the significant features of target objects. Research on these tailored loss functions is an active area of development.
@@ 6. RQ6
    Specific Datasets and preprocessing Techniques: Studies in the literature commonly use two publicly available datasets: FUSAR-Ship [17] and OpenSARShip [96]. Geometric preprocessing techniques, such as cropping, rotation, flipping and scaling, are frequently employed to improve the generalizability of ship classification models. However, the necessity for these techniques arises from the challenges associated with small and low-resolution vessel representations in the data. Super-resolution techniques offer a potential avenue for addressing these limitations [97, 98, 99, 89, 90].
@@ IV Background
Several techniques are employed for SAR ship classification using DL, as illustrated in Figure 2. This section is divided into four parts. The first part explores the various DL techniques used. The second part examines the impact of handcrafted features on the performance of DL models. As SAR images possess unique attributes, the third part discusses how these attributes are leveraged. Finally, the fourth part investigates the use of fine-tuning.
IV-A DL Techniques
@@
As apparent from Table II, we found the origin of SAR-based ship classification using DL in 2017 [20], when four different ship types detected in six VV-polarized COSMO-SkyMed scenes were classified using a simple CNN model with two convolution layers and two fully connected layers. The images of the detected targets were cropped to $64\times 64$ chips, four scenes were then used for training, one for testing and one for validation. The results showed that the model was able to learn the cargo ships better than the others, because of their high prevalence in the training set. The advancement of SAR platforms over time has resulted in images with better and better spatial resolutions. In [21], CNNs are used to treat images from TerraSAR-X, with a resolution of 3 meters per pixel. Four augmentation techniques are applied to form the training set: reflections (horizontal and vertical), rotations (-10 to 10 degrees), translation (-5 to 5 pixels) and noise addition (randomly changing pixel values to zero), achieving an F1 score of 94% by using multiple-resolution images.
@@
Two important aspects of a DL model are explored in [23]: an attention based CNN model and a novel loss function for class-imbalanced datasets, named “weighted distance measure”. A VH-polarized dataset with Ground Range Detected (GRD) images is divided into 4 classes, cargo, tanker, tug and others. The results show that the model is capable of learning class representation, achieving a recognition accuracy of 84%.
@@
In [27] a convolution network for detection and classification is evaluated. The input data are drawn from the PALSAR-2 L-band SAR with 10 m resolution in HV polarization. The results indicate that the CNN-based models can detect ships, but the classification performance is affected by ships that look alike.
@@
Ship classification can also be achieved using a sequence-based approach. In [29] a sequence input is constructed by capturing ten consecutive images of the same target at different times and rotation degrees. The authors employ a combination of CNN and long short-term memory (LSTM) with four categories of OpenSARShip data (boat, cargo, container, and oil tanker). Their findings suggest that the model learns improved class representations when using image sequences. Similarly, the authors of [74] first optimize the CNN design by reducing filter redundancy and network size, and then integrate LSTM layers to capture sequential dependencies and enhance discrimination. Comparative experiments against state-of-the-art CNNs (e.g., VGG-16, ResNet50, DenseNet121) show that the hybrid network achieves competitive accuracy while significantly reducing training time and complexity. These results highlight the potential of lightweight CNN–LSTM frameworks for efficient ship classification in resource-constrained settings.
DL Techniques for SAR Ship Classification Custom CNN Architectures [20, 21, 27, 23] Architecture enhancements [33, 75, 36, 57, 67, 62] Temporal / Hybrid (CNN–LSTM) [29, 74] Lightweight / NAS [84, 53, 56] Metric / Few-shot / Siamese [30, 60, 61, 51] Training strategies [50, 82, 64, 87] Data-centric (SR, domain) [90, 89, 65]
Figure 3: Overview of DL techniques applied to SAR ship classification. The methods are grouped into major categories, including core Custom CNN Architectures, architecture enhancements, temporal and hybrid approaches, lightweight and NAS-based models, metric and few-shot learning, training strategies, and data-centric methods. Representative studies for each category are indicated by citation keys.
@@
A CNN, when not fully connected, can be used as a feature extractor. The extracted features can then be fed to a classifier or further processed using methods like metric learning. Metric learning involves training a model to measure the similarity or distance between data points, often to cluster extracted features in a way that makes them distinguishable, thereby enhancing the model’s performance. In [30], the performance of a CNN-only approach is compared to a CNN combined with metric learning. The OpenSARship dataset is used, restricted to bulk carrier, container, and oil tanker classes, focusing on data with larger sample sizes. The features extracted using the CNN are then enhanced by concatenating additional features learned using metric learning. The results show that the CNN combined with metric learning outperforms the simple CNN.
@@
In [60], a novel Hierarchical Embedding Network with Center Calibration (HENC) is proposed for few-shot ship classification. HENC uses a CNN to extract hierarchical features at different scales, achieving an accuracy of 81.03% on a six-category OpenSARShip dataset (including cargo, tanker, tug, fishing, dredging and others) with few-shot training, thus demonstrating the effectiveness of deep models in learning class representations even with limited data.
@@
In [61], a deep metric learning approach is discussed, which leverages a pairwise representation fusion for multi-proxy learning and introduces a novel distribution loss to refine the sample distribution. This allows for both clustering similar classes together and training the model to make a clear distinction between all classes. The effectiveness of this three-pronged approach is evaluated on two optical ship datasets (FGSC-23 and FGSCR-42) and one SAR dataset (5-class FUSARShip).
@@
In [67], the authors introduce HDSS-Net, a novel DL architecture for ship classification in SAR images. This addresses challenges such as large variations in ship size, the elongated shapes of ships in SAR images, and intraclass diversity alongside interclass similarity. HDSS-Net employs a hierarchical design with three key modules. The first, a feature aggregation module (FAM), creates a multi-scale feature pyramid. The second, a Feature Boost module (FBM), captures the elongated features of ships through rectangular convolutions. Finally, a spherical space classifier (SSC) processes features in a spherical space to enhance class separability. Evaluations on OpenSARShip (three and six classes) and FUSAR-Ship (seven classes) demonstrate that HDSS-Net outperforms traditional CNN methods, highlighting its potential value for maritime surveillance applications.
@@
There exists a trade-off between spatial resolution and the cost of SAR imagery. Developing algorithms that achieve accurate results using medium-resolution data is essential for broader applications. DenseNet and metric learning, with a novel loss function incorporating deep metric learning [34], are used in [100] to perform SAR ship classification on three-class (tanker, container, bulk carrier) and five-class (tanker, container, bulk carrier, cargo, general cargo) datasets from OpenSARShip. To help the model learn accurate class representations, image triplets (an anchor image, a positive image, and a negative image) are used, and data augmentation techniques such as flipping, rotation, translation and Gaussian noise addition are applied to address overfitting. The results demonstrate superior performance in both the three-class and five-class experiments.
@@
Each CNN layer stores important information about features, but some information may be lost when features pass from one layer to another. In [33], a novel multi-scale CNN (MS-CNN) is proposed to utilize all the information from each convolutional layer. Using three classes from the OpenSARship dataset (bulk carrier, container ship and tanker), the MS-CNN, integrated with information from each convolutional layer, achieves a 5% better performance than the same CNN without integration. This improvement suggests that information loss occurs within the network.
@@
In [57], a novel deep network is proposed, with an inception-residual controller (IRC) module to improve the performance. The authors train the model on a five-category dataset from FUSARShip. The data undergo homogenization, normalization, and augmentation techniques such as flipping and rotation and, after 50 training epochs, the model with IRC achieves an accuracy of 98.71%, indicating that deep models trained on pre-processed data can significantly improve the classification performance.
Lightweight CNNs have been explored as means to reduce the computational burden of ship classification from high-resolution SAR data while maintaining competitive accuracy. A compact architecture called LN-SCNet is proposed in [84]. It integrates a preprocessing module to suppress speckle noise, dilated convolutions for initial feature extraction, and incorporates a feature fusion module that combines high-resolution features for small ships with deeper semantic features for large ships. To mitigate class imbalance, a seesaw loss is introduced, dynamically adjusting gradients across classes. Evaluated on 8 classes of the FUSAR-Ship dataset, LN-SCNet achieves 76.1% accuracy and an F1 score of 56.7%, outperforming other lightweight designs such as MobileNet and ShuffleNet while requiring only a fraction of their parameters and floating-point operations. These results suggest that carefully designed lightweight networks can offer an effective trade-off between efficiency and performance in SAR ship classification.
A plug-and-play Multi-scale Feature Spatial Attention Fusion (MFSAF) module is introduced in [75] to enhance CNN performance in SAR ship classification. This module integrates spatial attention with feature alignment, enabling CNNs to better capture ship features across multiple scales. Applied to several backbones (e.g., VGG-11, ResNet-18, DenseNet-121, EfficientNetV2, ConvNeXt), the MFSAF module consistently outperforms the baseline models, achieving an 81.9% accuracy on OpenSARShip2.0 when applied to ConvNeXt. These results highlight the adaptability of MFSAF and its effectiveness in improving spatial attention and feature fusion for robust ship classification in SAR imagery.
@@
A multi-scale dual-attention CNN that leverages information from convolutional layers is proposed in [36]. The authors employ a VGG-Net architecture for feature extraction and utilize a three-class dataset (cargo, container ship, and tanker) from OpenSARShip. The methodology involves modifying the VGG network to extract features from SAR images, followed by the application of squeeze-and-excitation (SE) blocks to focus on both channel-wise and positional context. SE blocks learn the importance of feature maps, promoting informative features and suppressing less useful channels. The results demonstrate the effectiveness of this method, achieving an accuracy of 92.66%. This success suggests the potential of attention mechanisms in this field. An enhanced vision transformer incorporating feature fusion modules to capture local information followed by SE attention is proposed in [62]. The effectiveness of this approach is evaluated on a three-class OpenSARship dataset, achieving an accuracy of 76.25%.
@@
In the training stage, the convergence of a DL model is controlled by several parameters, among which the learning rate is of paramount relevance. It determines how the model adjusts its internal parameters during training, and can either be fixed throughout the process or adjusted dynamically. Three learning rate management strategies and their impact on transfer learning are explored in [50]. The results suggest that a smooth learning rate change can improve the model’s ability to understand the data, leading to enhanced classification performance. The authors also investigate the effect of freezing different layers during transfer learning on a CNN, emphasizing the importance of layer analysis before fine-tuning. The models are first trained on the ten-category MSTAR dataset (SAR images of tanks, [101]) and then fine-tuned on a three-category OpenSARShip dataset (cargo, container ship, and bulk carrier). The results demonstrate that incorporating learning rate scheduling and thoughtful layer freezing significantly improves the model performance. This finding underscores the importance of employing effective training strategies for optimal SAR ship classification.
@@
Selecting optimal DL architectures for specific tasks can be challenging. Traditionally, researchers rely on exploring existing architectures and their applications. However, this approach often lacks a systematic methodology, leading to difficulties in finding the ideal architecture for a given classification problem. To address this challenge, a lightweight searched binary neural network (SBNN) is proposed in [53], obtained through network architecture search (NAS), quantized for weight reduction and incorporating patch shift processing to remove noise. The authors employ three (bulk carrier, container and tanker) and six-class (bulk carrier, container ship, tanker, cargo, general cargo and fishing) datasets from OpenSARShip, containing both VV and VH polarizations. They apply data augmentation methods such as random cropping and horizontal flipping. The SBNN achieves accuracy scores of 80.03% and 56.73% for three and six classes, respectively. NAS not only results in superior performance but also lead to a lightweight model. A novel mini-size searched convolutional metaformer (SCM) is introduced in [56], leveraging an updated NAS algorithm with progressive data augmentation (crop, flip, rotate, resize, cutout) to discover an efficient convolutional network baseline. This baseline is then enhanced with a transformer classifier to improve spatial awareness, and further augmented with a ConvFormer cell, combining the searched convolutional cell within a Metaformer block for superior feature extraction. Evaluated on VV- and VH-polarized datasets from OpenSARShip (3 categories, i.e. bulk carrier, container, and tanker) and FUSARShip (4 categories, i.e. bulk carrier, fishing, cargo, and tanker), SCM achieves accuracies of 82.06% and 63.9%, respectively, with a lightweight design (about half a million weights). In [51], a one-shot learning approach using a siamese network is proposed. The approach employs a 7-layer, fully connected feature fusion that enables the network to classify 16 categories within the OpenSARShip dataset. This methodology involves three stages: (i) preprocessing the images and converting them to three dimensions by adding two additional channels (actual, filtered, and inverted SAR Image), (ii) creating image triplets containing an anchor image, a positive image, and a negative image, and (iii) training the model using a loss function that ensures the anchor image to be closer to the positive image and farther from the negative image. The results demonstrate the model’s effectiveness for ship classification.
@@
The challenge of handling the long-tailed distribution problem in imbalanced SAR data is addressed in [64]. The authors propose a novel architecture that incorporates a multi-branch expert network with a dual-environment sampling strategy. This network consists of three classifiers trained to handle different distribution scenarios: long-tailed, balanced, and inverse long-tailed. Intraclass dual-environment sampling adjusts sampling rates within each class based on classification confidence, creating varied distributions. In [87], curriculum learning is used to address long-tailed distributions by introducing two curriculum methodologies for training, and also explores the importance of loss functions. Combined with center loss [102], this method enhances robustness and accuracy, outperforming existing techniques on the MSTAR and FUSAR datasets. Experiments confirmed significant improvements in recognizing both majority and minority classes, thus demonstrating the effectiveness of the proposed approach in balancing feature learning and improving overall and average accuracy.
Overfitting remains a key limitation in SAR ship classification, particularly in few-shot and noisy data scenarios. To address this, [82] presents a double reverse regularization strategy that leverages both offline and online self-distillation in a complementary manner. An Adaptive Weight Assignment (AWA) module is used to dynamically adjust the balance between the two sources of soft targets according to network performance, rather than relying on fixed schedules. Experiments on OpenSARShip and FUSAR-Ship show that this strategy consistently improves the performance of classical CNN backbones, achieving accuracy gains of up to 7% over baselines and surpassing other distillation-based regularization approaches. The results suggest that adaptive distillation can provide an effective path toward improved generalization under limited training data.
@@
Due to the resolution of SAR images, research on small-sized fishing vessels has not been given sufficient importance. However, as the resolution of SAR systems increases, there is a growing need to address the classification of fishing vessels. A novel architecture to this purpose is proposed in [65], consisting of three modules: (i) the multi-feature extraction (MUL) module, which extracts features using DenseBlocks and channel-wise concatenation, (ii) the feature fusion (FF) module, which aggregates features from different layers, and (iii) the parallel channel and spatial attention (PCSA) module, which applies channel and spatial attention. The paper also proposes a method for SAR and AIS integration to create a fishing vessel dataset. This methodology was tested on three fishing vessel classes, and the model achieved an accuracy of 89.7%.
The importance of super resolution in SAR ship classification is discussed in [90], whereas [89] emphasizes the importance of considering classification loss when enhancing the image resolutions of imbalanced datasets.